OpenStudy (anonymous):

x --- +33=34 80 What is X?

OpenStudy (mathhelp346):

X=80

OpenStudy (mathhelp346):

80/80=1 1+33=34

OpenStudy (anonymous):

x might b 1

OpenStudy (anonymous):

Consider the fact that any number over itself is 1. Then 80/80 is 1. 1 + 33 = 34, so you know that x = 80. In the same sense if this problem were written as. x/80 + 33 = 35, we would know that x = 160 as 160/80 = 2 Many people forget that a number divided by itself is 1. Happens to the best of us. Please Like Tutor Sean on facebook. It will help me help others, and I offer free help there with a community of other college tutors. Also check out http://www.tutorsean.net for free videos on math and related subjects.